the question is is how tall is this tree. and its a right triangle , the top is 30 ft and the bottom is 20 feet and the back of the triangle has an H how do i solve this
Answer by addingup(3677)

Do you mean like in the drawing below? If so,
H = sqrt(30^2-20^2) = sqrt(900-400) = sqrt(500) = 22.36